Many economists pointed out that in the 2008 recession V (velocity) fell
dramatically.
a. Why would V fall?
b. What would this do to Q or P (= MV, or V=PQ/M)?
c. Is your answer consistent with what actually happened?
d. Explain
1
Expert's answer
2016-10-26T13:08:08-0400
Many economists pointed out that in the 2008 recession V (velocity) fell dramatically.
a. As M*V = P*Q = Y, then velocity V would fall as the result of decrease in Y (GDP), because M (money supply) didn't change.
b. Q or P would also change, Q would decrease and P would increase.
c-d. So, my answer in a and b is consistent with what actually happened, because the price level P increased, the quantity of goods produced Q decreased, V decreased too.